# Explicit $7$-torsion in the Tate–Shafarevich groups of genus $2$ Jacobians

### v1

@@ Abstract.
We describe an algorithm which, on input a genus $2$ curve $C/\mathbb{Q}$ whose Jacobian $J/\mathbb{Q}$ has real multiplication by a quadratic order in which $7$ splits, outputs twists of the Klein quartic curve parametrising elliptic curves whose mod $7$ Galois representations are isomorphic to a sub-representation of the mod $7$ Galois representation attached to $J/\mathbb{Q}$. Applying this algorithm to genus $2$ curves of small conductor in families of Bending and Elkies–Kumar we exhibit a number of genus $2$ Jacobians whose Tate–Shafarevich groups (unconditionally) contain a non-trivial element of order $7$ which is visible in an abelian three-fold.
@@ 1. Introduction
Remark 1.2.
The genus $2$ Jacobian $J/\mathbb{Q}$ of conductor $3200^{2}$ in Table 1 was included in an appendix to [KS23] joint with Keller and Stoll where the strong Birch and Swinnerton-Dyer conjecture is also verified for $J$.
@@
$8$ $-10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
$8$ $165(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 39325 39325.c1
$37$ $-13(27x^{6}-54x^{5}-90x^{4}+228x^{3}+15x^{2}-90x-23)$ 73008 73008.n1
$8$ $-51(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 93925 93925.d1
$8$ $285(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 117325 117325.c1
$8$ $-46(x^{6}+6x^{5}-20x^{4}+240x^{3}+70x^{2}-84x+12)$ 203136 203136.i2
$8$ $-5(3x^{6}+12x^{5}+89x^{4}-56x^{3}-7x^{2}-132x+99)$ 211200 211200.c1
$8$ $465(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 216225 432450.ci1
$8$ $-30(11x^{6}-18x^{5}+47x^{4}+6x^{3}+71x^{2}+18x+27)$ 244800 244800.dc1
$8$ $-390(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 270400 270400.dc2
$8$ $-177(3x^{6}+12x^{4}-10x^{3}-12x+11)$ 281961 2819610*
$8$ $-22(x^{6}-24x^{5}+100x^{4}+102x^{3}-80x^{2}-132x-39)$ 302016 302016.p1
$8$ $-6(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 313920 313920.bb1
$8$ $-55(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 329725 659450*
$8$ $-110(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 387200 8905600*
$8$ $11(x^{6}+6x^{5}+11x^{4}-13x^{2}+6x-2)$ 423984 423984.by1
@@
The group $\Sha(A/K)$ is a torsion group and is conjectured to be finite. It is conjectured that for every prime number $p>0$ and each integer $g>0$ there exists an absolutely simple abelian variety (i.e., one which is not isogenous over $\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu$ to a product) of dimension $g$ for which $\Sha(A/\mathbb{Q})[p]\neq 0$. Indeed, it is even conjectured that $\Sha(A^{d}/\mathbb{Q})[p]\neq 0$ for a positive proportion of quadratic twists of a fixed abelian variety $A/\mathbb{Q}$ (see e.g., [BKLS21, Conjecture 1.1]).
In spite of this, for general values of $g$ and $p$, constructing an example of an absolutely simple $g$-dimensional abelian variety $A/\mathbb{Q}$ with an $p$-torsion element contained in $\Sha(A/\mathbb{Q})$ is an open problem. By allowing the dimension of $A$ to increase with $p$, Shnidman and Weiss [SW22] construct examples of absolutely simple abelian varieties with $\Sha(A/\mathbb{Q})[p]\neq 0$. Flynn and Shnidman [FS22] extended this result to show $\Sha(A/\mathbb{Q})[p]$ can be arbitrarily large.
@@
The central idea for proving Theorem 1.1 is to construct examples of genus $2$ curves $C/\mathbb{Q}$ with the property that there exists a $(7,\psi)$-congruence between an elliptic curve $E/\mathbb{Q}$ and the Jacobian $J=\operatorname{Jac}(C)$ of $C$, for some isogeny $\psi\colon J\to B$. Assuming that the local conditions are satisfied, it then suffices to show that $B(\mathbb{Q})/\psi J(\mathbb{Q})=0$ and that $E(\mathbb{Q})/7E(\mathbb{Q})\neq 0$ (which in practice is computationally less intensive than performing a $\psi$-descent on $J$). This approach is a mirror to that taken by Fisher [Fis16] who used it to visualise elements of order $7$ in the Tate–Shafarevich groups of elliptic curves. The main technical contribution of this article is construct examples of elliptic curves which are $(7,\psi)$-congruent to a genus $2$ Jacobian.
We ensure the existence of such an isogeny $\psi$ by choosing $J/\mathbb{Q}$ to have real multiplication (RM) by a real quadratic order $\mathcal{O}_{D}$ of fundamental discriminant $D>0$. Suppose that $7$ splits in $\mathcal{O}_{D}$ and we have an embedding $\mathcal{O}_{D}\subset\operatorname{End}_{\mathbb{Q}}(J)$. Writing $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ by abuse of notation we write $\mathfrak{p}\colon J\to B$ for the isogeny with kernel consisting of those $P\in J(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$ annihilated by $\mathfrak{p}$. In this case, $\ker\mathfrak{p}$ is isomorphic as a group to $(\mathbb{Z}/7\mathbb{Z})^{2}$ and under suitable hypotheses (see Lemma 2.2) comes equipped with a natural alternating bilinear pairing.
@@
Remark 1.4.
Since abelian surfaces $J/\mathbb{Q}$ with RM by $\mathcal{O}_{D}$ are modular (this follows from Serre’s conjecture [KW09a, KW09b]) we may associate to $J$ a weight $2$ newform with coefficients in $\mathcal{O}_{D}$ and level $\sqrt{N_{J}}$, where $N_{J}$ is the conductor of $J$ (in particular $N_{J}$ is a perfect square).
We compute these twists of $X(7)$ for examples of genus $2$ Jacobians of small conductor provided by the real multiplication families of Bending [Ben98, Ben99] and of Elkies–Kumar [EK14, CFM24]. By searching for rational points on these twists, we find a number of putative examples of $(7,\mathfrak{p})$-congruences between an elliptic curve $E/\mathbb{Q}$ and a genus $2$ Jacobian $J/\mathbb{Q}$. Adapting an approach of Fisher [Fis16, Section 6] we prove these congruences in 5.2.
For an abelian variety $A/\mathbb{Q}$ we write $A^{d}/\mathbb{Q}$ for the quadratic twist of $A$ by a squarefree integer $d\in\mathbb{Z}$. Note that simultaneous quadratic twists of $(7,\mathfrak{p})$-congruent pairs remain $(7,\mathfrak{p})$-congruent (cf. [FK22, Lemma 4.15]). To construct the examples in Theorem 1.1 we simply search for quadratic twists of the examples in 5.2 where there is a rank discrepancy between $E^{d}/\mathbb{Q}$ and $J^{d}/\mathbb{Q}$.
In addition to Theorem 1.1 we prove that there exist examples of such genus $2$ Jacobians with $7$-torsion in their Tate–Shafarevich groups and with real multiplication by $\mathcal{O}_{D}$ for several fundamental discriminants $D>0$.
Theorem 1.5.
For each $D=8$, $29$, $37$, $44$, and $57$ there exists an absolutely simple genus $2$ Jacobian $J/\mathbb{Q}$ with real multiplication by $\mathcal{O}_{D}$ such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$. Examples are furnished by the Jacobians of the curves $C\colon y^{2}=f(x)$ given in Table 2.
@@
Table 2. Examples of genus $2$ curves $C\colon y^{2}=f(x)$ whose Jacobians $J/\mathbb{Q}$ have real multiplication by the quadratic order $\mathcal{O}_{D}$ of discriminant $D$ and such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$. The subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides $7$ in $\mathcal{O}_{D}$. Explicit Weierstrass equations for the corresponding elliptic curves are given in [Fre]. The curves $C/\mathbb{Q}$ were generated from [Ben99, Theorem 4.1] and [EK14, CFM24]. Conductors were computed using [DD19].
In [KS23, A.3] it is observed that the Birch and Swinnerton–Dyer conjecture predicts that $\Sha(J^{-11}/\mathbb{Q})$ contains a subgroup isomoprhic to $(\mathbb{Z}/7\mathbb{Z})^{2}$, where $J/\mathbb{Q}$ is the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1. Since $J$ has RM by $\mathcal{O}_{8}$ it is natural to ask whether the $7$-torsion in $\Sha(J^{-11}/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence with an elliptic curve. By computing the relevant twists of $X(7)$ we give evidence that this is not the case (see Section 6).
Conjecture 1.6.
@@ 1.1. Outline of the paper
The outputs of the algorithm in Section 3.2 are not guaranteed to be correct, however in Section 4 we prove that the output is correct in many cases (for example for the curves in Tables 1 and 2). In particular, in Section 4 we prove that the twists we obtain are isomorphic to those which parametrise elliptic curves $(7,\mathfrak{p})$-congruent to $J$.
In Section 5 we prove Theorems 1.1 and 1.5 by proving that the pairs $(E,J)$ in Tables 1 and 2 are $(7,\mathfrak{p})$-congruent, and by checking that the local hypotheses in [Fis16, Theorem 2.2] are satisfied.
Finally in Section 6 we give explicit examples of the Klein quartic twists for the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1. By searching for rational points on these twists, we give evidence towards 1.6.
@@ 2. The modular curve $X(p)$ and its twists $X_{M}^{\pm}(p)$
[…]
@@ 2.1. The modular curve $X(p)$
The group $\Gamma\!_{p}$ of symplectic automorphisms of $\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$ acts naturally on $Y(p)$ on the right via $(E,\iota)\mapsto(E,\gamma\iota)$. As an abstract group $\Gamma\!_{p}$ is isomorphic to $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$, but it comes equipped with a non-trivial action of $\operatorname{Gal}(\mathbb{Q}(\mu_{p})/\mathbb{Q})$. The matrix ±$\pm I$ acts trivially on $Y(p)$ and therefore the action of $\Gamma\!_{p}$ factors through $\Gamma\!_{p}/\{\pm I\}$. This action extends to an action on $X(p)$, and the quotient realises the forgetful morphism $X(p)\to X(1)$ given by taking $j$-invariants.

### v2

@@ Abstract.
Let $C/\mathbb{Q}$ be a genus $2$ curve whose Jacobian $J/\mathbb{Q}$ has real multiplication by a quadratic order in which $7$ splits. We describe an algorithm which outputs twists of the Klein quartic curve which parametrise elliptic curves whose mod $7$ Galois representations are isomorphic to a sub-representation of the mod $7$ Galois representation attached to $J/\mathbb{Q}$. Applying this algorithm to genus $2$ curves of small conductor in families of Bending and Elkies–Kumar we exhibit a number of genus $2$ Jacobians whose Tate–Shafarevich groups (unconditionally) contain a non-trivial element of order $7$ which is visible in an abelian three-fold.
@@ 1. Introduction
Remark 2.
The genus $2$ Jacobian $J/\mathbb{Q}$ of conductor $3200^{2}$ in Table 1.1 was included in an appendix to [KS25] joint with Keller and Stoll where the strong Birch and Swinnerton-Dyer conjecture is also verified for $J$.
@@
88 $10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
88 $165(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 39325 39325.c1
$37$ $13(27x^{6}-54x^{5}-90x^{4}+228x^{3}+15x^{2}-90x-23)$ 73008 73008.n1
88 $51(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 93925 93925.d1
88 $285(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 117325 117325.c1
88 $62(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 184512 184512.bw1
88 $46(x^{6}+6x^{5}-20x^{4}+240x^{3}+70x^{2}-84x+12)$ 203136 203136.i2
88 $5(3x^{6}+12x^{5}+89x^{4}-56x^{3}-7x^{2}-132x+99)$ 211200 211200.c1
88 $465(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 216225 432450.ci1
88 $30(11x^{6}-18x^{5}+47x^{4}+6x^{3}+71x^{2}+18x+27)$ 244800 244800.dc1
88 $13(x^{6}-2x^{5}+3x^{4}x^{3}-7x^{2}-2x+1)$ 256880 51376.e1
88 $390(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 270400 270400.dc2
88 $177(3x^{6}+12x^{4}-10x^{3}-12x+11)$ 281961 2819610*
88 $22(x^{6}-24x^{5}+100x^{4}+102x^{3}-80x^{2}-132x-39)$ 302016 302016.p1
88 $6(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 313920 313920.bb1
88 $55(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 329725 659450*
88 $110(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 387200 8905600*
88 $11(x^{6}+6x^{5}+11x^{4}-13x^{2}+6x-2)$ 423984 423984.by1
Table $1.$ Examples of genus $2$ curves $C\colon y^{2}=f(x)$ whose Jacobians $J/\mathbb{Q}$ have conductor $N_{J}<(500\,000)^{2}$, and such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$ (we do not claim, nor expect, this list to be complete). The Jacobians have real multiplication by the quadratic order $\mathcal{O}_{D}$ of discriminant $D$ and the subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides 77 in $\mathcal{O}_{D}$. We write NN* for an elliptic curve of conductor NN which does not appear in the LMFDB (i.e., if N>500  [^0] N>500\,000). Explicit Weierstrass equations for the corresponding elliptic curves are given in [Fre]. The curves $C/\mathbb{Q}$ were generated from [Ben99, Theorem 4.1] and [EK14, CFM24]. Conductors were computed using [DD19].
@@ 88 $110(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 387200 8905600*
In spite of this, for general values of $g$ and $p$, constructing an example of an absolutely simple $g$-dimensional abelian variety $A/\mathbb{Q}$ with an $p$-torsion element contained in $\Sha(A/\mathbb{Q})$ is an open problem. By allowing the dimension of $A$ to increase with $p$, Shnidman and Weiss [SW22] construct examples of absolutely simple abelian varieties with $\Sha(A/\mathbb{Q})[p]\neq 0$. Flynn and Shnidman [FS25] extended this result to show $\Sha(A/\mathbb{Q})[p]$ can be arbitrarily large.
@@
The central idea for proving Theorem 1 is to construct examples of genus $2$ curves $C/\mathbb{Q}$ with the property that there exists a $(7,\psi)$-congruence between an elliptic curve $E/\mathbb{Q}$ and the Jacobian $J=\operatorname{Jac}(C)$ of $C$, for some isogeny $\psi\colon J\to B$. Assuming that the local conditions are satisfied, it then suffices to show that $B(\mathbb{Q})/\psi J(\mathbb{Q})=0$ and that $E(\mathbb{Q})/7E(\mathbb{Q})\neq 0$ (which in practice is computationally less intensive than performing a $\psi$-descent on $J$). This approach is a mirror to that taken by Fisher [Fis16] who used it to visualise elements of order $7$ in the Tate–Shafarevich groups of elliptic curves. The main technical contribution of this article is construct examples of elliptic curves which are $(7,\psi)$-congruent to a genus $2$ Jacobian.
We ensure the existence of such an isogeny $\psi$ by choosing $J/\mathbb{Q}$ to have real multiplication (RM) by a real quadratic order $\mathcal{O}_{D}$ of fundamental discriminant $D>0$. Suppose that $7$ splits in $\mathcal{O}_{D}$ and we have an embedding $\mathcal{O}_{D}\subset\operatorname{End}_{\mathbb{Q}}(J)$. Writing $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ by abuse of notation we write $\mathfrak{p}\colon J\to B$ for the isogeny with kernel consisting of those $P\in J(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$ annihilated by $\mathfrak{p}$. In this case, $\ker\mathfrak{p}$ is isomorphic as a group to $(\mathbb{Z}/7\mathbb{Z})^{2}$ and under suitable hypotheses (see Lemma 8) comes equipped with a natural alternating bilinear pairing.
@@
Remark 4.
Since abelian surfaces $J/\mathbb{Q}$ with RM by $\mathcal{O}_{D}$ are modular (this follows from Serre’s conjecture [KW09a, KW09b]) we may associate to $J$ a weight 22 newform with coefficients in $\mathcal{O}_{D}$ and level $\sqrt{N_{J}}$, where $N_{J}$ is the conductor of $J$ (in particular $N_{J}$ is a perfect square).
We compute these twists of $X(7)$ for examples of genus $2$ Jacobians of small conductor provided by the real multiplication families of Bending [Ben98, Ben99] and of Elkies–Kumar [EK14, CFM24]. By searching for rational points on these twists, we find a number of putative examples of $(7,\mathfrak{p})$-congruences between an elliptic curve $E/\mathbb{Q}$ and a genus $2$ Jacobian $J/\mathbb{Q}$. Adapting an approach of Fisher [Fis16, Section 6] we prove these congruences in 22.
For an abelian variety $A/\mathbb{Q}$ we write $A^{d}/\mathbb{Q}$ for the quadratic twist of $A$ by a squarefree integer $d\in\mathbb{Z}$. Note that simultaneous quadratic twists of $(7,\mathfrak{p})$-congruent pairs remain $(7,\mathfrak{p})$-congruent (cf. [FK22, Lemma 4.15]). To construct the examples in Theorem 1 we simply search for quadratic twists of the examples in 22 where there is a rank discrepancy between $E^{d}/\mathbb{Q}$ and $J^{d}/\mathbb{Q}$.
In addition to Theorem 1 we prove that there exist examples of such genus $2$ Jacobians with $7$-torsion in their Tate–Shafarevich groups and with real multiplication by $\mathcal{O}_{D}$ for several fundamental discriminants $D>0$.
Theorem 5.
For each $D=8$, $29$, $37$, $44$, and $57$ there exists an absolutely simple genus $2$ Jacobian $J/\mathbb{Q}$ with real multiplication by $\mathcal{O}_{D}$ such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$. Examples are furnished by the Jacobians of the curves $C\colon y^{2}=f(x)$ given in Table 1.2.
@@
Table 1.2. Examples of genus $2$ curves $C\colon y^{2}=f(x)$ whose Jacobians $J/\mathbb{Q}$ have real multiplication by the quadratic order $\mathcal{O}_{D}$ of discriminant $D$ and such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$. The subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides 77 in $\mathcal{O}_{D}$. We write NN* for an elliptic curve of conductor NN which does not appear in the LMFDB (i.e., if N>500  [^0] N>500\,000). Explicit Weierstrass equations for the corresponding elliptic curves are given in [Fre]. The curves $C/\mathbb{Q}$ were generated from [Ben99, Theorem 4.1] and [EK14, CFM24]. Conductors were computed using [DD19].
In [KS25, A.3] it is observed that the Birch and Swinnerton–Dyer conjecture predicts that $\Sha(J^{-11}/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$, where $J/\mathbb{Q}$ is the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1. Since $J$ has RM by $\mathcal{O}_{8}$ it is natural to ask whether the $7$-torsion in $\Sha(J^{-11}/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence with an elliptic curve. By computing the relevant twists of $X(7)$ we give evidence that this is not the case (see Section 6).
Conjecture 6.
@@ 1.1. Outline of the paper
The outputs of the algorithm in Section 3.2 are not guaranteed to be correct, however in Section 4 we prove that the output is correct in many cases (for example for the curves in Tables 1.1 and 1.2). In particular, in Section 4 we prove that the twists we obtain are isomorphic to those which parametrise elliptic curves $(7,\mathfrak{p})$-congruent to $J$.
In Section 5 we prove Theorems 1 and 5 by proving that the pairs $(E,J)$ in Tables 1.1 and 1.2 are $(7,\mathfrak{p})$-congruent, and by checking that the local hypotheses in [Fis16, Theorem 2.2] are satisfied.
Finally in Section 6 we give explicit examples of the Klein quartic twists for the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1. By searching for rational points on these twists, we give evidence towards 6.
@@ 2. The modular curve $X(p)$ and its twists $X_{M}^{\pm}(p)$
Let $K$ be a field of characteristic zero. A symplectic abelian group over $K$ is a pair $(M,e_{M})$ where $M$ is a $G_{K}$-module equipped with a ($G_{K}$-equivariant) alternating, bilinear pairing $e_{M}\colon M\times M\to\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu^{\times}$. We equip $\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$ with the natural alternating pairing $\langle(\zeta,n),(\xi,m)\rangle=\zeta^{m}\xi^{-n}$.
@@ 2.1. The modular curve $X(p)$
The group $\Gamma\!_{p}$ of symplectic automorphisms of $\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$ acts naturally on $Y(p)$ on the right via $(E,\iota)\mapsto(E,\gamma\iota)$. As an abstract group $\Gamma\!_{p}$ is isomorphic to $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$, but it comes equipped with a non-trivial action of $\operatorname{Gal}(\mathbb{Q}(\mu_{p})/\mathbb{Q})$. The matrix $\pm I$ acts trivially on $Y(p)$ and therefore the action of $\Gamma\!_{p}$ factors through $\Gamma\!_{p}/\{\pm I\}$. This action extends to an action on $X(p)$, and the quotient realises the forgetful morphism $X(p)\to X(1)$ given by taking $j$-invariants.